OpenStudy (ell):

Use the properties of exponents to simplify each expression. (-7x^2) x (-4x^5)

OpenStudy (anonymous):

Multiply the coefficents and add the exponents :D

OpenStudy (ivettef365):

just multiply (-7) * (-4) then (x^2)(x^5) to multiply exponents you have to add x^a * x^b = X^(a+b)

OpenStudy (campbell_st):

use the process of Numbers Letters Powers Signs so \[-7 \times -4\times x^{2 + 5}\] the index law for multiplication of the same base is add the powers.

OpenStudy (ell):

So it would be 28^7 ?

OpenStudy (campbell_st):

well you have missed the letter or base... the number and power are correct

OpenStudy (ell):

28x^7 ?

OpenStudy (campbell_st):

thats correct

OpenStudy (anonymous):

Yes @Ell :)

OpenStudy (ell):

Thank you so much guys!! GBU!! (: <3

OpenStudy (anonymous):

Your always welcome :)

OpenStudy (kenljw):

(-7x^2) x (-4x^5) If the x above means multiplication if should be expressed (-7x^2)X(-4x^5) or (-7x^2)(-4x^5) This eliminates possible confusion, generally the second is preferred like xy procedure factor out coefficients (-7)(-4)(x^2)x^5) use principles -- = +, (x^n)(x^m) = x^(n+m)

OpenStudy (ell):

Thanx! it really helps! :)